The equation x^2+y^2+z^2+8x−6y+2z+17=0 represents a sphere. The radius of the sphere is,
Select one:
a. √3

b. 9
c. 3
d. 17
e. √17
1
Expert's answer
2020-05-13T19:49:18-0400

Consider the equation of circle "x^2+y^2+z^2+8x-6y+2z+17=0"


Complete square for each coordinate points "x,y" and "z" as,


"x^2+8x+y^2-6y+z^2+2z+17=0"


"x^2+2(4)x+4^2-4^2+y^2-2(3)y+3^2-3^2+z^2+2(1)z+1^2-1^2+17=0"


"(x+4)^2-16+(y-3)^2-9+(z+1)^2-1+17=0"


"(x+4)^2+(y-3)^2+(z+1)^2-26+17=0"


"(x+4)^2+(y-3)^2+(z+1)^2-9=0"


"(x+4)^2+(y-3)^2+(z+1)^2=9"


"(x+4)^2+(y-3)^2+(z+1)^2=3^2"


Therefore, the radius of the sphere is "3" .


Hence, option (c) is correct.